Status Report

NASA ISS On-Orbit Status 2 May 2009

By SpaceRef Editor
May 2, 2009
Filed under , , ,
NASA ISS On-Orbit Status 2 May 2009
NASA ISS On-Orbit Status 2 May 2009

All ISS systems continue to function nominally, except those noted previously or below. Saturday – half-duty day for the crew. Space Day Family Day!

FE-1 Barratt continued his first session of sleep logging for the experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) from his Actiwatch to the HRF-1 (Human Research Facility 1) laptop as part of a week-long run. This is similar to Barratt’s BCD (Baseline Data Collection) which was performed pre-flight for comparison. [To monitor the crewmember’s sleep/wake patterns and light exposure, Dr. Mike wears a special Actiwatch device which measures the light levels encountered by them as well as his patterns of sleep and activity throughout the Expedition and uses the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days.]

FE-2 Wakata began the day by closing the CBEF (Cell Biology Experiment Facility) door in the Saibo Rack to provide room in the JPM (JEM Pressurized Module). Later in the day, the CBEF door was opened again to reduce the humidity level in the incubator.

The crew performed the regular weekly three-hour task of thorough station cleaning. [“Uborka”, usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE’s sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

As part of the house cleaning, Padalka conducted regular maintenance inspection & cleaning on fan screens in the FGB (TsV2), DC1 (V3), and SM (VPkhO, VPrK, FS5, FS6 & FS9).

The CDR also did the periodic maintenance of the Russian POTOK-150MK (150 micron) air purification subsystems in the SM and FGB by cleaning the pre-filters with a vacuum cleaner with narrow nozzle attachment and later restarting the POTOK in automatic mode.

After shutdown of the EarthKAM (EK/Earth Knowledge Acquired by Middle School Students) experiment down by ground commanding, Michael Barratt dismantled the setup in Node-2 and stowed the equipment.

For his VolSci (Voluntary Science) program session today, the FE-1 configured the BCAT (Binary Colloid Alloy Test) payload in the Node-2, checked it out and started the experiment. At the end of the experiment run, Barratt moved the A31p laptop from the BCAT location to the BISE experiment location in the Lab for the next experiment. [Mike set up the experiment along with camcorder and DCS-760 still camera, finding and photographing crystals of different colors in samples 8, 9 and 10, then setting up the camera for taking multiple flash angle images of sample 4, followed by homogenizing sample 4 and taking test photos. The FE-1 then initiated the activity of automated photography of sample 4 with the Kodak DCS-760 digital still camera controlled by the EarthKAM software running on an SSC-13 (Station Support Computer 13) for about 87 days.]

FE-2 Wakata’s VolSci program for today was another session with the BISE (Bodies in the Space Environment) experiment, complete with video coverage, investigating the relative contributions of internal and external cues to self-orientation during and after zero-G exposure. After setting up the camcorder for recording the activity, configuring the “Neurospat” hardware and activating the BISE software on its A31p SSC (Station Support Computer) laptop, Wakata then had ~90 min for completing the experiment protocol. [The CSA (Canadian Space Agency)-sponsored BISE experiment studies how astronauts perceive Up and Down in microgravity. The specific objective of the BISE project is to conduct experiments during long-duration microgravity conditions to better understand how humans first adapt to microgravity and then re-adapt to normal gravity conditions upon return to earth. This experiment involves comparisons of preflight, flight, and post-flight perceptions and mental imagery, with special reference to spaceflight-related decreases in the vertical component of percepts. The test involves having subjects view a computer screen through a cylinder that blocks all other visual information. The astronauts will be presented with background images with different orientations relative to their bodies.]

In Kibo, FE-2 Wakata prepared his camera equipment at window JPM1F8 for the JAXA EPO (Educational Program Operation) PHOTO-MOON, then took images of the moon from that window for ~10 minutes. [The purpose of the MOON EPO is to create a musical score using moon photos taken from Kibo while the crewmember is floating naturally in micro-G. The position of the moon in the photos is regarded as the sound pitch (quarter note, eighth note, etc.), e.g., when the moon is at the top of the picture, it is a high pitch. The project will emphasize the difference between the photo (and score) taken from the Earth and taken from Kibo. Photos taken with the 400 mm lens use the moon craters to create the musical score. Crater size determines sound length. At least 80 photos are necessary each with the 200mm and 400mm lens.]

Mike performed the daily procedure of flushing [the PWD (Potable Water Dispenser) ambient line with ~50mL of water (into a towel/Ziploc bag). PWD water is currently cleared only for hygienic use, pending Russian approval.

CDR Padalka completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

The crew completed their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (CDR, FE-1, FE-2), ARED advanced resistive exercise device (FE-1, FE-2) and VELO bike with bungee cord load trainer (CDR).

Afterwards, Wakata downloaded the exercise data file to the MEC (Medical Equipment Computer) for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on ARED, followed by their erasure on the HRM storage medium (done six times a week).

At ~10:00am EDT, the crew conducted their regular WPC (Weekly Planning Conference) with the ground, discussing next week’s “Look-Ahead Plan” (prepared jointly by MCC-Houston and TsUP-Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

At ~11:18am, Barratt powered up the SM’s amateur radio equipment (Kenwood VHF transceiver with manual frequency selection, headset, & power supply) and at ~11:23am conducted a ham radio session with visitors assembled at the National Air & Space Museum (NASM) in Washington, DC. This public exchange, mostly with students selected from the audience, was part of the Space Day Family Day events celebrating NASA’s 51st anniversary and 40th year anniversary of the first moon landing. [Check out ]

Weekly Science Update (Expedition Nineteen — Week 4

3-D SPACE: Koichi completed his third session of 3D-SPACE. Thank you for performing this experiment.

AgCam (Agricultural Camera): “Mike, after you enabled power to AgCam, ground teams performed additional power-up testing of the payload. AgCam is both sending health and status data and receiving and processing commands, but can neither read certain internal status parameters nor power on subsystem components. Both of these conditions indicate there is a problem with communication between the laptop parallel port and the AgCam electronics. Since you already confirmed the cable is securely connected, we are developing further on-orbit troubleshooting. This may take a week to develop to check if the laptop is properly configured and operating correctly. If this can resolve the problem, final steps for AgCam setup and activation can be completed in short order, and science imagery can begin.”

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

ALTEA DOSI (NASA/ASI): Standing by.

BCAT-3/4 (Binary Colloidal Alloy Test 3/4): Nothing new.

Biological Rhythms (JAXA): Ongoing. Two runs are for Barratt, two are planned for De Winne.

BISE (CSA, Bodies in the Space Environment): “The data acquired during BISE third session as voluntary science was very good. A 4th session is planned on May 2nd to cover the adaptation phase to microgravity. Next sessions will be scheduled at the end of your flight. Thanks again, Mike, for the excellent work.”


CARDIOCOG-2: Complete.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Ongoing.

CFE (Capillary Flow Experiment): Reserve.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Complete.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CIR (Combustion Integrated Rack): Nothing new.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

Commercial 2 (JAXA): Cosmo Flower mission was completed by Wakata (slipped from I-18)

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.

DomeGene (JAXA): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): The 5-days 5-hours of operations have yielded over 1200 images commanded by 93 schools/classes involving nearly 7,000 students. The EarthKAM library of images from 31 sessions onboard the ISS stands at over 31,650.

EDR (European Drawer Rack): The rack is continuously active in support of the PCDF (Protein Crystallization Diagnostic Facility) experiment. Two anomalies are currently under investigation.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

ENose (Electronic Nose): Operating.

EPM (European Physiology Module): Underway.

EPO (Educational Payload Operations): Reserve.

EPO J-Astro Report (JAXA): “The reports are very well received. Reading audience is enjoying the reports.”

EPO Space Clothes (JAXA): Space Clothes activity was completed by Wakata (slipped from I-18).

EPO Hiten (Dance, JAXA): Planned task, now completely finished. PI expressed satisfaction for the results. We appreciate Koichi’s great job. We have finished all the tasks initially planned for Increment 18.

EPO Moon Score (JAXA): “Photo Moon #2 is complete. Now we are downloading the data. Next session is scheduled on Saturday. We appreciate your work last weekend.

EPO Try Zero-G (JAXA): First part of Try Zero-G was performed on 4/27. Koichi’s performance

was perfect and showed successfully good examples of the difference between 0G and 1G. JAXA had press tour to show his performance with downlink video, then 5 TV and 10 newspaper reporters attended and all of them enjoyed watching his performance with his message. Next session is planned for next week.

EPO Kibo Kids Tour (JAXA): “Recorded video was screened on TKSC Open house event on Sunday and was received well. Number of visitors was almost doubled than last year. Thank you for your work.”

EPO Spiral Top (JAXA): Planned task was completely finished. PI expressed satisfaction for the results. We appreciate Koichi’s great job. We have finished all the tasks initially planned for Increment 18.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): Platform nominal. DOSTEL, EXPOSE, FIPEX, MEDET: On-going science acquisition – nominal; DEBIE-2: the instrument continues to generate empty science packets at regular intervals (of 30 to 34hrs). Science acquisition is pursued with regular power cycling of the instrument (work-around)

FACET (JAXA): “Beautiful facet crystal growth is observed. We will finish the experiments using Cell1 after 8 more experiments. We are planning a crew task to rotate the Experiment Cell to Start with Cell #2 which contains different composition of solution to explore more science.”

FSL (Fluid Science Laboratory): Activated for MMA measurements in support of PCDF.

GEOFLOW: No report.

HDTV System (JAXA): To be launched by HTV1.

Holter (JAXA): One of two sessions for FE-2 is slipped from Increment 18. Scheduled on 5/21.

HQPC (JAXA): To be launched by 34P.


IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): Complete.

Integrated Immune: Complete.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Completed.

Marangoni Experiment for ISS in JAXA FPEF (Fluid Physics Experiment Facility): In progress.

MAXI (JAXA): Exposed Payload, to be launched on 2JA.

Micro-G Clay (JAXA EPO): Complete.

MISSE (Materials ISS Experiment): Ongoing.

Moon Photography from ISS (JAXA EPO): One run performed on the last day of Increment 18.

MSG-SAME (Microgravity Science Glovebox): Complete.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.


PADLES (JAXA, Area PADLES 3; Passive Dosimeter for Lifescience Experiment in Space): Continuing radiation dose accumulation.

PCDF-PU (Protein Crystallization Diagnostic Facility – Process Unit): Growth cycle EP4 Cycle3, ended on Friday, 4/24. EP1 Cycle2 was started on 4/24. Over the weekend EP4 was left at +25degC and images were taken before the start of EP4 Cycle4. EP4 Cycle4 was continued after the PCDF reboot on 4/27 and restarted after both EDR anomalies. The current cycle has been running since 4/29.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.


RadGene & LOH (JAXA): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SEDA-AP (JAXA): Exposed Payload,- to be launched by 2JA.

SHERE (Shear History Extensional Rheology Experiment): Complete.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Mike, you are currently in process of completing your first week of sleep logging. We are going to be targeting every third week for scheduling your sleep logging. Thanks for continuing to support this experiment.

SMILES (JAXA): Exposed Payload, to be launched by HTV1.

SOLAR (Solar Monitoring Observatory): Acquired science until Sun observation window ended on 4/29.

SOLO (Sodium Loading in Microgravity): No report.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): Reserve.

SPICE (Smoke Point In Co-flow Experiment): Restarted SPICE operations on GMT 4/29.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


VLE (Video Lessons ESA): VLE-1 completed.

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete/Planned (2J/A Stage).

CEO (Crew Earth Observations): “Through 4/82, the ground has received a total of 4,598 of ISS CEO imagery for review and cataloguing.”

No CEO (Crew Earth Observation) photo targets uplinked for today.

CEO photography can be studied at this “Gateway” website: (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this morning, 8:27am EDT [= epoch])

Mean altitude — 351.6 km

Apogee height – 357.7 km

Perigee height — 345.5 km

Period — 91.58 min.

Inclination (to Equator) — 51.64 deg

Eccentricity — 0.0000806

Solar Beta Angle — 37.8 deg (magnitude increasing)

Orbits per 24-hr. day — 15.72

Mean altitude loss in the last 24 hours — 49 m

Revolutions since FGB/Zarya launch (Nov. 98) — 59876

Significant Events Ahead (all dates Eastern Time, some changes possible!):

05/06/09 — Progress M01M/32P undocking & deorbit

05/07/09 — Progress M-02M/33P launch (on Soyuz-U, 51st rocket of this type)

05/11/09 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)

05/12/09 — Progress M-02M/33P docking

05/27/09 — Soyuz TMA-15/19S launch

05/29/09 — Soyuz TMA-15/19S docking (FGB nadir)

Six-person crew on ISS

06/05/09 — Russian EVA-22

06/10/09 — Russian EVA-23

06/13/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD

07/17/09 – Progress M-02M/33P undock & deorbit

07/20/09 — Soyuz TMA-14/18S relocation (from SM aft to DC1)

07/24/09 — Progress 34P launch

07/26/09 — Progress 34P docking (SM aft)

08/06/09 — STS-128/Discovery/17A – MPLM (P), LMC

09/01/09 — H-IIB (JAXA HTV-1) launch – tentative

09/07/09 — H-IIB (JAXA HTV-1) berth

09/30/09 — Soyuz TMA-16/20S launch

10/02/09 — Soyuz TMA-16/20S docking (SM aft, until MRM2 w/new port)

10/08/09 — H-IIB (JAXA HTV-1) unberth

10/11/09 – Soyuz TMA-14/18S undock

10/15/09 — Progress 35P launch

11/10/09 — 5R/MRM2 (Russian Mini Research Module, MIM2) on Proton — tentative

11/12/09 — STS-129/Atlantis/ULF3 – ELC1, ELC2

12/07/09 — Soyuz TMA-17/21S launch

12/26/09 — Progress 36P launch

02/03/10 — Progress 37P launch

02/??/10 — STS-130/Endeavour/20A – Node-3 + Cupola — tentative

02/11/10 — STS-131/Atlantis/19A – MPLM(P), LMC — tentative

04/02/10 — Soyuz TMA-18/22S launch

04/08/10 — STS-132/Discovery/ULF4 – ICC-VLD, MRM1 — tentative

04/27/10 — Progress 38P launch

05/29/10 — Soyuz TMA-19/23S launch

05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4 — tentative

06/??/10 – ATV2 – Ariane 5 (ESA)

06/25/10 — Progress 39P launch

08/11/10 — Progress 40P launch

09/29/10 — Soyuz TMA-20/24S launch

12/??/11 — Proton 3R/MLM w/ERA.

10/19/10 — Progress 41P launch

12/??/11 – 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton.

SpaceRef staff editor.